# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v2

@@ VI-C3 MENN of a Triangle
The Steiner inellipse which is tangent to the edges of the triangle at their midpoints, is the MENN of the triangle [49]. As shown in Fig. 5, we denote the 2-D Cartesian coordinates of the vertices of the triangle as $v^{*}=(v^{*}_{0},v^{*}_{1})^{\mathrm{T}},*=\{\textrm{i},\textrm{ii},\textrm{iii}\}$. The center $v^{o}$ and two conjugate diameters $f_{1},f_{2}\in\mathbb{R}^{2}$ of the Steiner inellipse can be written as
@@
TABLE I: Success Rate of Different Methods for Generating Convex Polytopes Containing Seed across Different Scenarios and Seed Types
[…]
Point Seed Line Seed Polytope Seed
   FIRI IRIS[1] Galaxy[9] RILS[2]    FIRI IRIS[1] Galaxy[9] RILS[2]    FIRI IRIS[1] Galaxy[9] RILS[2]
2-D Sparse 100 98.4 100 100 100 74.7 81.6 100 100 88.9 95.8 97.0
Medium 100 97.0 100 100 100 53.9 64.2 100 100 79.7 91.1 95.3
Dense 100 96.6 100 100 100 48.5 39.1 100 100 69.7 73.9 90.6
3-D Sparse 100 99.1 100 100 100 96.5 79.1 100 100 96.2 85.1 98.0
Medium 100 97.8 100 100 100 78.2 61.3 100 100 70.6 78.0 88.3
Dense 100 96.2 100 100 100 59.6 47.0 100 100 37.3 45.1 70.6
@@ VII Evaluation and Benchmark
To comprehensively demonstrate the outstanding advantages of the proposed algorithm, FIRI, in terms of efficiency, quality, and manageability, we design extensive benchmarks comparing FIRI against several state-of-the-art convex polytope generation algorithms. Additionally, the efficiency of FIRI lies in the methods that we develop specifically for the two optimization problems (Line 1 and 1 of Algorithm 1) within FIRI. To demonstrate their effectiveness, we conduct comparative experiments for these methods which are proposed for solving minimum-norm with small dimension but massive constraints and for solving MVIE, respectively. All benchmarks are conducted on an Intel Core i7-12700KF CPU under Linux, and are implemented in C++11 without any explicit hardware acceleration.
VII-A Comparison of Generating Free Convex Polytope
Based on the description and analysis of several state-of-the-art algorithms for generating free convex polytopes in Sec. II, we benchmark the proposed algorithm FIRI against IRIS [1], Galaxy [9] and RILS [2]. Since Gao’s method [8] relies on modeling the environment as a grid map and can only operate in near real-time with low map resolution, it is not compared here. Additionally, Galaxy can be regarded as an enhancement of Savin’s method [12], both of which are based on space inversion, thus we directly adopt Galaxy for benchmark. For IRIS and our proposed method FIRI, we set the same parameter $\rho=0.02$, that is, the stopping condition of them is that the volume of the MVIE of the convex hull obtained in this iteration grows less than $2\%$ from the last iteration. For Galaxy and RILS, we used the default parameters.
To begin with, we compare the capability of these methods to adapt to different types of seed and obstacle inputs. As evidenced by the results in Tab. II, FIRI demonstrates the highest level of adaptability to various inputs. Subsequently, based on the reported adaptability, we conduct extensive benchmarks to validate the superior performance of FIRI.
VII-A1 Manageability
We conduct benchmark in a complex environment of $50\times 50~{}m$ size (with a height of $10m$ for the 3-D case), where we generate random obstacles by using Perlin noise [55]. For each test, we randomly generate a collision-free seed in the environment as input. The boundary of each convex hull generation algorithm is constrained to be a square (cube for 3-D) with side length $6~{}m$ centered on the seed’s center and parallel to the coordinate axis. And the obstacle input is the points within the boundary of the square in the map. When the seed input is a line, we set its length to $1.5~{}m$. When the seed input is a convex polytope, we set it as a $1.5\times 0.75\times 0.25~{}m$ rectangular for the 3-D case, or a $1.5\times 0.75~{}m$ rectangle for the 2-D case. We generate the experimental environments with different obstacle densities, whose numbers of input obstacles are shown in Tab. IV. For each density, we create different $10$ environments. Furthermore, for each environment, we conduct $500$ random trials for each type of seed input.
TABLE II: Comparison of Adaptability of Different Methods for Generating Free Convex Polytope
@@ VII-A1 Manageability
To be fair, we perform several adjustments in different seed cases based on Tab. II. When the seed is a line, for Galaxy and IRIS, which can only use one point as the seed input, we use both the endpoints and midpoints of the line as seed inputs for them to compute $3$ convex polytopes. We then select the convex hull with the highest degree of line containment as the result. When the seed is a convex polytope, FIRI can directly take the polytope as input. IRIS and Galaxy take each vertex and the center point of the seed as inputs to compute multiple corresponding convex hulls. For RILS, which is adapted to use a line as input, we compute multiple convex hulls using the seed’s diagonals as inputs. For each of the above three methods, we choose the convex hull that maximizes the inclusion of the seed as the respective result.
TABLE IV: Input Obstacles Number of Different Scenarios
Scenario Input Obstacle Number
avg std min max
2-D Sparse 246.7 111.0 44 525
Medium 1157.6 277.8 553 1683
Dense 3007.5 515.2 1443 4048
3-D Sparse 453.6 119.6 303 670
Medium 2677.8 613.6 1524 3929
Dense 12659.0 764.9 11032 14536
@@
We calculate the success rate of each method in generating a convex polytope that fully contains the seed input, as shown in Tab. I. In addition, for a more intuitive presentation, we present an example of the results of each method in an application of whole-body planning that has a need for manageability. As shown in Fig. 11(b), we use a rectangular robot as the seed, and the details of the application will be described in Sec. VIII-A1. Due to greedy approach of IRIS in seeking the largest possible volume of the convex hull, it performs poorly in terms of manageability, to the extent that it cannot guarantee to contain the seed even when the seed is just a single point. While RILS can ensure to contain the line seed, it is not sufficient to guarantee that the generated convex hull contains the seed when the seed is represented as a convex hull. As for Galaxy, its heuristic cut method does not ensure that the seed is included either, when the seed is a line or a convex polytope. In contrast, benefiting from RsI, FIRI ensures manageability for all three types of seed inputs across various obstacle densities.
VII-A2 Efficiency and Quality
We conduct experiments to compare efficiency and quality in the aforementioned random environments. The obstacle input and boundary settings remain the same as in Sec. VII-A1. We record the computation time for each algorithm when seeded with a point as input. If IRIS generates a convex polytope during the iteration process that does not include the seed, we force IRIS to terminate prematurely and return the polytope from the previous iteration that includes the seed as the result. Since both RILS and FIRI guarantee manageability for line seeds, we document their respective time overhead when the seed input is a line as well. In addition, FIRI maintains its manageability when only a single iteration is performed, we also record the outcomes of a single iteration of FIRI, denoted as FIRI(SI).
@@ VII-A2 Efficiency and Quality
To provide an intuitive comparison of the size of convex polytopes generated by each algorithm, we take the proposed FIRI, which aims to maximize the convex hull volume, as the baseline. Specifically, we report the ratio of the volume of the convex hull obtained by each algorithm to the volume of the one obtained by FIRI, as demonstrated in Fig. 6. The efficiency results are presented in Tab. III, where, for clarity, we distinguish between iterative and non-iterative methods using dashed lines in the case of point seed. In addition, since the number of input obstacles is the primary factor influencing the computational efficiency, we showcase the the quantity of input obstacles in different density environments in Tab. IV to provide readers with an insight into the computational efficiency of each algorithm. As the results illustrated in Fig. 6 and Tab. III, both IRIS and FIRI iteratively compute larger convex polytopes by continuously inflating the MVIE, resulting in similar sizes. However, in IRIS, the SDP-based MVIE solving method consumes significant computation time [1], whereas FIRI achieves significant efficiency improvements in MVIE calculations as shown in Sec. VII-C, leading to remarkably higher efficiency compared to IRIS. Moreover, FIRI achieves a computational time that is within three times the time of non-iterative RILS that does not involve MVIE computations. The other three non-iterative methods yield smaller convex polytopes, but achieve better computational efficiency than IRIS in Tab. III. Galaxy, however, is even less efficient than FIRI due to the fact that the Quickhull [30] algorithm involved degrades in the scene to which Galaxy corresponds, which is particularly noticeable in the 3D case. As for FIRI(SI), similar to RILS, it directly updates the convex hull and can generate polytope of comparable size to RILS, but with higher computational efficiency. Moreover, benefiting from the manageability brought about by the RsI, if the objective is to obtain a free convex polytope that includes the seed as fast as possible, without pursuing maximum volume, we believe that FIRI(SI) is a more suitable choice.
@@
Apart from using point representation for obstacles as used above, employing convex polytopes is also a commonly used method for representing the environment. Among the the above methods shown in Tab. II,, only FIRI and IRIS support obstacle inputs in the form of convex hulls, thus we further compare these two algorithms. For each test we generate a certain number of randomly distributed obstacles in a closed space and select the center of the space as the seed input. This closed space serves as the boundary input to each method. We conducted benchmarks by setting the number $N_{v}$ of the vertices of a single convex obstacle to vary from $3$ to $10^{3}$. Compared to the case where obstacles are represented by points, the relative volumes of obstacles that are convex polytopes are larger, thus we vary the number of obstacles to vary from $10$ to $10^{3}$, which we denote as spare, medium and dense in the results shown in Fig. 7. For each the number of the vertices of each different number of obstacles, we perform $500$ tests. The increasing number of obstacles, the number of halfspaces generated for convex hull construction tends to stabilize, thus the computational efficiency advantage provided by the porposed MVIE solution gradually diminishes. However, benefiting from the efficient SDMN we propose for halfspace computation, FIRI can still maintain a significant advantage over IRIS, even in scenarios with dense obstacles, with a reduction of over 95% in computational requirements.
@@ VII-A3 Case Study of Quality
In order to demonstrate the significance of convex hull quality and its impact on trajectory planning, we conduct an experiment on trajectory planning based on different convex hull generation methods in the random environment as shown in Fig. 8(a). We first generate a collision-free path connecting the start and end points using RRT* [56]. The obtained path can be viewed as a set of connected line segments, based on which we generate convex hulls subsequently. As indicated in Tab. I, only FIRI and RILS exhibit manageability over line seed. Thus we build corridors for comparison based on these two methods, which correspond to iterative and non-iterative strategies, regarding the pursuit of maximum volume, respectively. The process of generating a safe flight corridor is as follows: We sequentially traverse each line segment on the path. If the line segment is already included in the previously generated polytopes, we move on to the next segment. Otherwise, we use this line as a seed to generate a new convex polytope. Due to manageability, the generated set of convex polytopes must have an intersection between two neighboring pairs, forming a corridor as the result. Based on the generated corridor, we adopt GPOPS-II [57] to obtain the optimal trajectory constrained within the corridor. This collocation-based method transcribes the trajectory optimization problem into a constrained Nonlinear Programming using the Gauss pseudospectral method, which is then solved by the well-established NLP solver SNOPT [58]. In GPOPS-II, each trajectory phase is confined within one polytope, and we set the feasibility constraints for velocity and acceleration as $3m/s$ and $6m/s^{2}$, and set the time weight as $20$. As depicted in Fig. 8(b), RILS generates narrow convex polytopes in the area marked by the black dashed circles, resulting in limited maneuvering space for trajectory optimization. Consequently, the trajectory constrained in the corridor generated by RILS exhibits a conservative behavior in the marked area . In contrast, FIRI, due to its pursuit of maximizing convex hulls, is capable of generating larger corridors, providing greater spatial flexibility in trajectory planning.
For the minimum-norm (16), we compare SDMN which is proposed in Sec. IV with several cutting-edge general-purpose QP solvers: the parametric active-set algorithm qpOASES [14], the operator-splitting based first-order approximation algorithm OSQP [15] and the interior-point method-based quadratic programming approximation algorithm HPIPM [16]. We use the open source code  [^1] https://github.com/coin-or/qpOASES [^2] https://github.com/osqp/osqp [^3] https://github.com/giaf/hpipm of these solvers to obtain the solution of (16) respectively. For qpOASES and HPIPM, we use the default parameters. However, we find that the default parameters of OSQP are insufficient when the constraint size becomes large. Therefore, we adjust the relative convergence tolerance to $10^{-12}$ and the maximal number of iterations to $10^{6}$, denoted as OSQP, while the implementation with the default parameters are denoted as OSQP default. As for the proposed SDMN, we do not require any additional parameter settings.
$\psi_{N}=\mathcal{L}_{\psi}({Ey^{*}-f}),$ (61)
where $y^{*}$ is the corresponding solution of each method, $E$ and $f$ are the parameters used to define the constraints of the minimum-norm (16), and $\mathcal{L}_{\psi}(\cdot)$ denotes a function that takes the absolute value of the largest element of the input vector. The precision $\psi_{N}$ indicates the degree of proximity of the solution of each method to the most active constraints of the input minimum-norm. The performance is reported in Fig. 9 and Tab. V. The results demonstrate that SDMN exhibits remarkably high computational efficiency, surpassing other methods by orders of magnitude. Additionally, SDMN provides analytical solutions, achieving high accuracy without the need for additional parameters.
@@ VII-B Comparison of Solving Strictly Convex Small Dimensional Minimum-Norm with Massive Constraints
TABLE V: Comparison of Precisions $\psi_{N}$ between Different Methods for Solving Small Dimensional Minimum-Norm

### v3

@@ VI-C3 MENN of a Triangle
The Steiner inellipse tangent to the edges of the triangle at their midpoints, is the MENN [52]. As shown in Fig. 5, we denote the 2-D Cartesian coordinates of the vertices of the triangle as $v^{*}=(v^{*}_{0},v^{*}_{1})^{\mathrm{T}},*=\{\textrm{i},\textrm{ii},\textrm{iii}\}$. The center $v^{o}$ and two conjugate diameters $f_{1},f_{2}\in\mathbb{R}^{2}$ of the Steiner inellipse can be written as
@@ VII Evaluation and Benchmark
To comprehensively demonstrate the outstanding advantages of FIRI, in terms of efficiency, quality, and manageability, we design extensive benchmarks comparing FIRI against several state-of-the-art convex polytope generation algorithms. Additionally, the efficiency of FIRI lies in the methods that we develop specifically for the two optimizations (Line 1 and 1 of Algorithm 1). To demonstrate their effectiveness, we conduct comparative experiments for these methods which are proposed for solving minimum-norm with small dimension but massive constraints and for solving MVIE, respectively. All benchmarks run on Intel Core i7-12700KF CPU, Linux, using C++14 without hardware acceleration.
TABLE I: Comparison of Adaptability of Different Methods for Generating Free Convex Polytope
@@
TABLE II: Input Obstacles Number of Different Scenarios
Scenario Input Obstacle Number
avg std min max
2-D Sparse 246.7 111.0 44 525
Medium 1157.6 277.8 553 1683
Dense 3007.5 515.2 1443 4048
3-D Sparse 453.6 119.6 303 670
Medium 2677.8 613.6 1524 3929
Dense 12659.0 764.9 11032 14536
VII-A Comparison of Generating Free Convex Polytope
Based on the description and analysis of several state-of-the-art algorithms for generating free convex polytopes in Sec. II, we benchmark the proposed FIRI against IRIS [1], Galaxy [9] and RILS [2]. Since Gao’s method [8] relies on modeling the environment as a grid map and can only operate in near real-time with low map resolution, it is not compared here. Additionally, Galaxy can be regarded as an enhancement of Savin’s method [12], both of which are based on space inversion, thus we directly adopt Galaxy for benchmark. For IRIS and our proposed method FIRI, we set the same parameter $\rho=0.02$, that is, the stopping condition of them is that the volume of the MVIE of the convex hull obtained in this iteration grows less than $2\%$ from the last iteration. For Galaxy and RILS, we used the default parameters.
To begin with, we compare the capability of these methods to adapt to different types of seed and obstacle inputs. As evidenced by the results in Tab. I, FIRI demonstrates the highest level of adaptability to various inputs. In Tab. II, the term of adaptability of polytope-type obstacles denotes the ability to handle polytope obstacles directly without discretization. As polytope representation is a widely adopted approach for environmental modeling [58, 59], the capability to process polytope obstacles directly is worthwhile. Subsequently, based on the reported adaptability, we conduct extensive benchmarks to validate the superior performance of FIRI.
VII-A1 Manageability
We conduct benchmark in a complex environment of $50\times 50~{}m$ size (with a height of $10m$ for the 3-D case), where we generate random obstacles by using Perlin noise [60]. For each test, we randomly generate a collision-free seed in the environment as input. The boundary of each convex hull generation algorithm is constrained to be a square (cube for 3-D) with side length $6~{}m$ centered on the seed’s center and parallel to the coordinate axis. And the obstacle input is the points within the boundary of the square in the map. When the seed input is a line, we set its length to $1.5~{}m$. When the seed input is a convex polytope, we set it as a $1.5\times 0.75\times 0.25~{}m$ rectangular for the 3-D case, or a $1.5\times 0.75~{}m$ rectangle for the 2-D case. We generate the experimental environments with different obstacle densities, whose numbers of input obstacles are shown in Tab. II. For each density, we create different $10$ environments. Furthermore, for each environment, we conduct $500$ random trials for each type of seed input.
To be fair, we perform several adjustments in different seed cases based on Tab. I. When the seed is a line, for Galaxy and IRIS, which can only use one point as the seed input, we use both the endpoints and midpoints of the line as seed inputs for them to compute $3$ convex polytopes. We then select the convex hull with the highest degree of line containment as the result. When the seed is a convex polytope, FIRI can directly take the polytope as input. IRIS and Galaxy take each vertex and the center point of the seed as inputs to compute multiple corresponding convex hulls. For RILS, which is adapted to use a line as input, we compute multiple convex hulls using the seed’s diagonals as inputs. For each of the above three methods, we choose the convex hull that maximizes the inclusion of the seed as the respective result.
We calculate the success rate in generating a convex polytope that fully contains the seed, as shown in Tab. LABEL:tab:success_rate_of_manageability. In addition, for a more intuitive presentation, we present an example of the results of each method in an application of whole-body planning that has a need for manageability. As shown in Fig. 12(b), we use a rectangular robot as the seed, and the details of the application will be described in Sec. VIII-A2. Due to greedy approach of IRIS in seeking the largest possible volume of the convex hull, it performs poorly in terms of manageability, to the extent that it cannot guarantee to contain the seed even when the seed is just a single point. While RILS can ensure to contain the line seed, it is not sufficient to guarantee that the generated convex hull contains the seed when the seed is represented as a convex hull. As for Galaxy, its heuristic cut method does not ensure that the seed is included either, when the seed is a line or a convex polytope. In contrast, although FIRI employs a similar approach to IRIS and RILS in computing halfspaces through ellipsoid inflation for polytope generation, the integration of our proposed RsI provides a distinct advantage: each halfspace computed by FIRI is guaranteed to contain the seed. This fundamental property ensures that the resulting polytope invariably encompasses the seed, thus maintaining manageability.
@@ VII-A1 Manageability
VII-A2 Efficiency and Quality
We conduct experiments to compare efficiency and quality in the aforementioned random environments. using the same settings of obstacles and boundary as Sec. VII-A1. We record the computation time for each algorithm with point seed. If IRIS generates a convex polytope during the iteration process that does not include the seed, we force IRIS to terminate prematurely and return the polytope from the previous iteration that includes the seed as the result. Since both RILS and FIRI guarantee manageability for line seeds, we document their respective time overhead when the seed input is a line as well. In addition, since the non-iterative method RILS essentially represents a single iteration of IRIS without MVIE computation, we also record the outcomes of a single iteration of FIRI, denoted as FIRI(SI).
@@ VII-A2 Efficiency and Quality
To provide an intuitive comparison of the size of convex polytopes generated by each algorithm, we take the proposed FIRI, which aims to maximize the convex hull volume, as the baseline. Specifically, we report the ratio of the volume of the convex hull obtained by each algorithm to the volume of the one obtained by FIRI, as demonstrated in Fig. 6. The efficiency results are presented in Tab. III, where, for clarity, we distinguish between iterative and non-iterative methods using dashed lines in the case of point seed. In addition, since the number of input obstacles is the primary factor influencing the computational efficiency, we showcase the quantity of input obstacles in different density environments in Tab. II to provide readers with an insight into the computational efficiency of each algorithm. As the results illustrated in Fig. 6, both IRIS and FIRI iteratively compute larger convex polytopes by continuously inflating the MVIE, resulting in similar sizes. However, in IRIS, the SDP-based MVIE solving method consumes significant computation time [1], whereas FIRI achieves significant efficiency improvements in MVIE calculations as shown in Sec. VII-C, leading to remarkably higher efficiency compared to IRIS as shown in Tab. III. Moreover, FIRI achieves a computational time that is within three times the time of non-iterative RILS without involving MVIE. The other three non-iterative methods yield smaller polytopes, but achieve better efficiency than IRIS in Tab. III. Galaxy, however, is even less efficient than FIRI due to the fact that Quickhull [33] involved degrades in the scene to which Galaxy corresponds, which is particularly noticeable in the 3D case. As for FIRI(SI), similar to RILS, it directly updates the convex hull and can generate polytope of comparable size to RILS, but with higher computational efficiency. Moreover, benefiting from the manageability brought about by the RsI, if the objective is to obtain a free convex polytope that includes the seed as fast as possible, without pursuing maximum volume, we believe that FIRI(SI) is a more suitable choice.
@@
Apart from using point representation for obstacles as used above, employing convex polytopes is also a commonly used method for representing the environment. Although methods such as RILS can handle polytope obstacles through discretizing them into points, this approach may incur increased computational burden when discretizing large-scale simple convex polytope, and certain discretization methods (like gridding) may compromise environmental modeling accuracy. Hence, we consider the capability to directly process polytope obstacles to be valuable. Among above methods in Tab. I, only FIRI and IRIS support obstacle inputs in the form of convex hulls, thus we further compare these two algorithms. For each test we generate a certain number of randomly distributed obstacles in a closed space and select the center of the space as the seed input. This closed space serves as the boundary. We conduct benchmarks by setting the number $N_{v}$ of the vertices of one obstacle to vary from $3$ to $10^{3}$. Compared to the case using point obstacles, the relative volumes of obstacles that are polytopes are larger, thus we vary the number of obstacles to vary from $10$ to $10^{3}$, which we denote as spare, medium and dense in the results shown in Fig. 7. For each the number of the vertices of each different number of obstacles, we perform $500$ tests. The increasing number of obstacles, the number of halfspaces generated for convex hull construction tends to stabilize, thus the computational efficiency advantage provided by the porposed MVIE solution gradually diminishes. However, benefiting from the efficient SDMN we propose for halfspace computation, FIRI can still maintain a significant advantage over IRIS, even in scenarios with dense obstacles, with a reduction of over 95% in computational requirements.
@@ VII-A3 Case Study of Quality
To demonstrate the significance of convex hull quality and its impact on trajectory planning, we conduct an experiment on trajectory planning based on different convex hull generation methods in the random environment as shown in Fig. 8(a). Using RRT*[61], we generate a collision-free path between start and end points. This path can be viewed as a set of connected line segments, based on which we generate convex hulls subsequently. As indicated in Tab. LABEL:tab:success_rate_of_manageability, only FIRI and RILS exhibit manageability over line seed. Thus we build corridors for comparison based on these two methods, which correspond to iterative and non-iterative strategies, regarding the pursuit of maximum volume, respectively. The process of generating a safe flight corridor is as follows: We sequentially traverse each line segment on the path. If the line segment is already included in the previously generated polytopes, we move on to the next segment. Otherwise, we use this line as a seed to generate a new convex polytope. Due to manageability, the generated set of convex polytopes must have an intersection between two neighboring pairs, forming a corridor. Based on the generated corridor, we adopt GPOPS-II [62] to obtain the optimal trajectory constrained within the corridor. This collocation-based method transcribes the trajectory optimization into a constrained Nonlinear Programming using the Gauss pseudospectral method, which is then solved by the well-established NLP solver SNOPT [63]. In GPOPS-II, each trajectory phase is confined within one polytope, and we set the feasibility constraints for velocity and acceleration as $3m/s$ and $6m/s^{2}$, and time weight as $20$. As depicted in Fig. 8(b), RILS generates narrow polytopes in the area marked by black dashed circles, resulting in limited maneuvering space for trajectory optimization. Consequently, the trajectory constrained in the corridor generated by RILS exhibits a conservative behavior in the marked area . In contrast, FIRI, due to its pursuit of maximizing convex hulls, is capable of generating larger corridors, providing greater spatial flexibility in trajectory planning.
@@
TABLE IV: Comparison of Precisions $\psi_{N}$ between Different Methods for Solving Small Dimensional Minimum-Norm
@@ VII-B Comparison of Solving Strictly Convex Small Dimensional Minimum-Norm with Massive Constraints
For minimum-norm (16), we compare SDMN with several cutting-edge general-purpose QP solvers: the parametric active-set algorithm qpOASES [14], the operator-splitting based first-order approximation algorithm OSQP [15] and the interior-point method-based quadratic programming approximation algorithm HPIPM [16]. We use the open source code  [^2] https://github.com/coin-or/qpOASES [^3] https://github.com/osqp/osqp [^4] https://github.com/giaf/hpipm of these solvers to obtain the solution of (16) respectively. For qpOASES and HPIPM, we use the default parameters. However, we find that the default parameters of OSQP are insufficient when the constraint size becomes large. Therefore, we adjust the relative convergence tolerance to $10^{-12}$ and the maximal number of iterations to $10^{6}$, denoted as OSQP, while the implementation with the default parameters are denoted as OSQP default. As for the proposed SDMN, we do not require any additional parameter settings.
$\psi_{N}=\mathcal{L}_{\psi}({Ey^{*}-f}),$ (61)
where $y^{*}$ is the corresponding solution of each method, $E$ and $f$ are the parameters used to define the constraints of the minimum-norm (16), and $\mathcal{L}_{\psi}(\cdot)$ denotes a function that takes the absolute value of the largest element of the input vector. The precision $\psi_{N}$ indicates the degree of proximity of the solution of each method to the most active constraints of the input minimum-norm. The performance is reported in Fig. 9 and Tab. IV. The results demonstrate that SDMN exhibits remarkably high computational efficiency, surpassing other methods by orders of magnitude. Additionally, SDMN provides analytical solutions, achieving high accuracy without the need for additional parameters.